Growth​ Company's current share price is $19.85 and it is expected to pay a $0.85 dividend per share next year.
After​ that, the​ firm's dividends are expected to grow at a rate of 3.7% per year.

a. What is an estimate of Growth​ Company's cost of​ equity?

b. Growth Company also has preferred stock outstanding that pays a $1.90 per share fixed dividend.
If this stock is currently priced at $28.10​, what is Growth​ Company's cost of preferred​ stock?

c. Growth Company has existing debt issued three years ago with a coupon rate of 6.3%.
The firm just issued new debt at par with a coupon rate of 6.4%.
What is Growth​ Company's cost of​ debt?

d. Growth Company has 4.7 million common shares outstanding and 1.2 million preferred shares​ outstanding,
and its equity has a total book value of $50.2 million.
Its liabilities have a market value of $20.1 million.
If Growth​ Company's common and preferred shares are priced as in parts (a​) and (b​),
what is the market value of Growth​ Company's assets?

e. Growth Company faces a 38% tax rate.
Given the information in parts (a​) through (d​),
and your answers to those​ problems, what is Growth​ Company's WACC?

​Note: Assume that the firm will always be able to utilize its full interest tax shield.
Answer #1
a) Cost of Company
= (D1 / P0) + g
Where,
D1 = Dividend of Next Year = $0.85
P0 = Current Share Price = $19.85
g = Expected Growth Rate = 3.7% = 0.037
So,
Cost of Company
= (D1 / P0) + g
= ($0.85 / $19.85) + 0.037
= 0.0428 + 0.037
= 0.0798
i.e. 7.98%
b) Cost of Preferred Stock
= Fixed Dividend / Current Price
= $1.90 / $28.10
= 0.0676
i.e. 6.76%
c) The Pre Tax Cost of Debt is the YTM of the Company on Current Debt.
Growth Company just issued new debt at par with Coupon Rate of 6.4%.
So Coupon Rate will be the pre Tax Cost of Debt i.e. 6.4%
Pre Tax Cost of Debt = 6.40%
After Tax Cost of Debt
= Pre Tax Cost of Debt*(1-Tax Rate)
= 6.40%*(1-38%)
= 6.40%*0.62
= 3.968%
d) Market Value of Company's Assets
= Marekt Value Debt + Market Value of Preferred Stock + Market Value of Common Shares
= Marekt Value Liabilities + (No. of Shares*Current Stock Price) + (No. of Shares*Current Stock Price of Common Shares)
= $20,100,000 + (1200000*$28.10) + (4700000*$19.85)
= $20,100,000 + $33,720,000+ $93,295,000
= $147,115,000
e) Market Weight Formula = Market Value of respective Source / Total Market Value
Source Market Value Market Weight Formula Market Weight Cost Cost*Weight
Equity $93,295,000 = 93295000 / 147115000               0.6342 7.98% 5.06%
Preferred Stock $33,720,000 = 33720000 / 147115000               0.2292 6.76% 1.55%
Debt $20,100,000 = 20100000 / 147115000               0.1366 3.968% 0.54%
$147,115,000 7.15%
WACC = 7.15%